OpenStudy (anonymous):

Choose the point-slope form of the equation below that represents the line that passes through the point (6, -3) and has a slope of 1/2 x - 2y = 12 y + 3 = (x - 6) y = x - 6 y - 6 = (x + 3)

OpenStudy (smilie2015):

y-y1=m(x-x1) y-6 =1/2(x+3) y-6= 1/2x+3/2 y-1/2x=3/2+6 y-1/2x =\[\frac{{3 + 12}}{2}\] y-1/2x =15/2 if u want to simplify more than multiply by 2 on both sides 2y-x=15

OpenStudy (smilie2015):

i think in this way u will solve

OpenStudy (anonymous):

so the answer would be the third one?